OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

@ali2x2

OpenStudy (ali2x2):

Let D = b^2 − 4ac This is known as the discriminant

OpenStudy (ali2x2):

If the discriminant D is a perfect square, then ax^2 + bx + c can be factored

OpenStudy (ali2x2):

Compare 2x^2 + 7x + 3 to ax^2 + bx + c

OpenStudy (ali2x2):

what are the values of 'a', 'b' and 'c'?

OpenStudy (ali2x2):

a = 2, b = 7 and c = 3

OpenStudy (ali2x2):

So D = b^2 - 4ac D = 7^2 - 4(2)(3) ... Plug in a = 2, b = 7 and c = 3 D = 49 - 24 D = 25

OpenStudy (ali2x2):

is 25 a perfect square?

OpenStudy (ali2x2):

yes or no?

OpenStudy (anonymous):

yes

OpenStudy (ali2x2):

so 2x^2 + 7x + 3 can be factored

OpenStudy (ali2x2):

then the ending value is 25

OpenStudy (ali2x2):

they just want the value of b^2 - 4ac

OpenStudy (ali2x2):

you type in 25

OpenStudy (ali2x2):

that's the final value of b^2 - 4ac

OpenStudy (ali2x2):

where a = 2, b = 7 and c = 3

OpenStudy (anonymous):

is 25 the final answerr

OpenStudy (ali2x2):

yes thats the final value of b^2 - 4ac where a = 2, b = 7 and c = 3

OpenStudy (ali2x2):

final value : final value of b^2 - 4ac = 25 where :a = 2, b = 7 and c = 3